# Modeling Innovation Ecosystem Dynamics through Interacting Reinforced Bernoulli Processes

### v1

@@ Appendix B Proofs of the main results and technical details for the statistical test
Using (B.3) and the relation $\zeta_{t+1}(x)/\zeta_{t}(x)=1+O(1/t)$, we obtain
(B.6) ${\mathbf{B}}_{t+1}$ $={\mathbf{B}}_{t}-\frac{1}{t+1}\left(\gamma^{*}{\mathbf{B}}_{t}-{\mathbf{A}}_{t}\right)$
$\qquad+\frac{1}{t+1}\Delta{\mathbf{M}}_{t+1}+O(\zeta_{t}(1-\gamma^{*})/t^{2}){\boldsymbol{1}}\,,$
where again $O(\zeta_{t}(1-\gamma^{*})/t^{2})=O(1/t^{1+\gamma^{*}})$. Finally, we observe that, by (B.2), we have
$t^{-(1-\gamma^{*})}E[\Delta{\mathbf{M}}_{t+1}\Delta{\mathbf{M}}_{t+1}^{\top}\,|\mathcal{F}_{t}]$ $\stackrel{{\scriptstyle a.s.}}{{\sim}}t^{1-\gamma^{*}}diag\left(P_{t,1}(1-P_{1,t}),\dots,P_{t,N}(1-P_{t,N})\right)$
$\stackrel{{\scriptstyle a.s.}}{{\longrightarrow}}\begin{cases}diag({\mathbf{P}}_{\infty})\quad&\mbox{for }\gamma^{*}<1\\ diag(P_{\infty,1}(1-P_{\infty,1}),\dots,P_{\infty,N}(1-P_{\infty,N}))\quad&\mbox{for }\gamma^{*}=1\;(\mbox{and so }{\mathbf{u}}={\boldsymbol{1}})\end{cases}$
=$\quad=\begin{cases}\widetilde{P}_{\infty}diag({\mathbf{u}})\quad&\mbox{for }\gamma^{*}<1\\ \widetilde{P}_{\infty}(1-\widetilde{P}_{\infty})I\quad&\mbox{for }\gamma^{*}=1\;(\mbox{and so }{\mathbf{u}}={\boldsymbol{1}})\end{cases}$
and we recall that, as proven before, $\widetilde{A}_{t}={\mathbf{v}}^{\top}{\mathbf{A}}_{t}=\zeta_{t}(1-\gamma^{*}){\mathbf{v}}^{\top}{\mathbf{P}}_{t}\stackrel{{\scriptstyle a.s.}}{{\to}}\widetilde{P}_{\infty}$. Therefore, the pair $({\mathbf{A}}_{t},{\mathbf{B}}_{t})_{t}$ satisfies the dynamics and the conditions required in the general central limit theorem proven in (Aletti et al. 2025, Appendix A), provided we assume ${\mathcal{R}e}(\gamma^{*}_{2})/\gamma^{*}<1/2$, where $\gamma^{*}_{2}$ is an eigenvalue of $\Gamma$ different from $\gamma^{*}$ with highest real part, that is $\gamma^{*}_{2}\in Sp(\Gamma)\setminus\{\gamma^{*}\}$ with ${\mathcal{R}e}(\gamma^{*}_{2})=\max\{{\mathcal{R}e}(\gamma):\gamma\in Sp(\Gamma)\setminus\{\gamma^{*}\}\}$. Hence, we can apply the statistical tools based on that result and described in (Aletti et al. 2025, Appendix B and C). In particular, in the mean-field case we have ${\mathbf{v}}=N^{-1}{\boldsymbol{1}}$, ${\mathbf{u}}={\boldsymbol{1}}$ and $\gamma^{*}_{2}=\gamma^{*}(1-\iota)$ so that, when $\iota>1/2$, we have
$(2\iota-1)t^{\gamma^{*}}\frac{\|{\mathbf{B}}_{t}-{\mathbf{v}}^{\top}{\mathbf{B}}_{t}\|^{2}}{{\mathbf{v}}^{\top}{\mathbf{B}}_{t}}\stackrel{{\scriptstyle d}}{{\longrightarrow}}\chi^{2}(N-1).$
B.1. General results
@@
(B.7) ${\mathbf{A}}_{t+1}={\mathbf{A}}_{t}-\frac{1}{t+1}(\phi^{*}I-\Phi^{\top}){\mathbf{A}}_{t}+\frac{1}{t+1}\Phi^{\top}\Delta{{\mathbf{M}}}_{t+1}+{\mathbf{R}}_{A,t+1}\$
@@
[…]
Let ${\mathbf{u}}$ and ${\mathbf{v}}$ be the left and the right eigenvectors of $\Phi$ associated to $\phi^{*}$, with strictly positive components and such that ${\mathbf{v}}^{\top}{\boldsymbol{1}}=1\quad\mbox{and}\quad{\mathbf{v}}^{\top}{\mathbf{u}}=1.$ (Recall that $\phi^{*}$ is real and simple and it is possible to choose the components of these vectors all strictly positive because of the Frobenius-Perron theory). Set $\widetilde{A}_{t}={\mathbf{v}}^{\top}{\mathbf{A}}_{t}$.
@@
Proof. By multiplying equation (B.7) by ${\mathbf{v}}^{\top}$ we obtain
(B.8) $\widetilde{A}_{t+1}=\widetilde{A}_{t}+\frac{1}{t+1}\phi^{*}\Delta\widetilde{M}_{t+1}+{\mathbf{v}}^{\top}{\mathbf{R}}_{A,t+1}\,,$
@@
(B.9) $\widetilde{A}_{t+1}-\widetilde{A}_{0}=\sum_{n=0}^{t}(\widetilde{A}_{n+1}-\widetilde{A}_{n})=\phi^{*}\breve{M}_{t+1}+\sum_{n=0}^{t}\widetilde{R}_{A,n+1}\,.$
Hence, since $\sup_{t}E[\,|\widetilde{A}_{t}|\,]<+\infty$ (by (iii)) and $\sum_{t}1/t^{1+\beta}<+\infty$, we also have $\sup_{t}E[\,|\breve{M}_{t}|\,]<+\infty$. Therefore, $(\breve{M}_{t})_{t}$ is a martingale bounded in $L^{1}$ and so it converges almost surely to an integrable random variable $\breve{M}_{\infty}$. It follows from (B.9) that $(\widetilde{A}_{t})$ converges almost surely to an integrable random variable $\widetilde{A}_{\infty}$. Moreover, we obtain
(B.10) $(\widetilde{A}_{t}-\widetilde{A}_{\infty})^{2}$ $=\big{(}\phi^{*}(\breve{M}_{t}-\breve{M}_{\infty})+\sum_{n\geq t}\widetilde{R}_{A,n+1}\big{)}^{2}$
@@
Now, we are going to prove that, under assumption (iv), we have $\sup_{t}(\widetilde{A}_{t})^{2}<+\infty$ so that (by (B.9)) $(\breve{M}_{t})_{t}$ is a martingale bounded in $L^{2}$ and so $\breve{M}_{\infty}$ is square-integrable and $\breve{M}_{t}$ converges in quadratic mean to it. By (B.10) this fact obviously implies that $\widetilde{A}_{t}$ converges in quadratic mean to $\widetilde{A}_{\infty}$.
@@
Therefore, from (B.8), since $\sup_{t}E[\widetilde{A}_{t}]<+\infty$, we get
@@
where we have used (iv) in order to say that the first series is finite. Therefore, assuming ${\mathbf{A}}_{0}$ (and so $\widetilde{A}_{0}$) square-integrable, we have $\sup_{t}E[(\widetilde{A}_{t})^{2}]<+\infty$. ∎∎
Remark B.1 (non-negative case).
In particular, condition (iii) is verified when $\widetilde{A}_{t}$ is non-negative for each $t$. Indeed, if $\widetilde{A}_{t}$ is non-negative, we have, for each $t$,
@@
and thus, since the last series is finite and $\widetilde{A}_{0}$ integrable, we have $\sup_{t}E[\widetilde{A}_{t}]<+\infty$, that is (iii) is verified.
@@
Proof. We firstly want to prove that we can neglect the term $R_{A,t+1}$ in the dynamics (B.7) of ${\mathbf{A}}_{t}$.
@@
(B.11) $V^{\top}{\mathbf{u}}=U^{\top}{\mathbf{v}}=0,\quad V^{\top}U=U^{\top}V=I,\quad I={\mathbf{u}}{\mathbf{v}}^{\top}+UV^{\top}\,.$
@@
(B.12) ${\mathbf{A}}_{t+1}$ $=\Big{(}I-\frac{1}{t+1}U(I\phi^{*}-D)V^{\top}\Big{)}{\mathbf{A}}_{t}$
@@
(B.13) ${\mathbf{A}}_{t+1}$ $=\ C_{m_{0},t}{\mathbf{A}}_{m_{0}}+\sum_{k=m_{0}}^{t}\frac{1}{k+1}C_{k+1,t}\Phi^{\top}\Delta{\mathbf{M}}_{k+1}$
@@
Moreover, setting for any $x\in{\mathbb{C}}$ with ${\mathcal{R}e}(x)/(m_{0}+1)<1$, $p_{m_{0}}(x)=1$ and $p_{k}(x)=\prod_{m=m_{0}}^{k}(1-\tfrac{x}{m+1})$ for $k\geq m_{0}$ and $F_{k+1,t}=\frac{p_{t}(x)}{p_{k}(x)}$ for $m_{0}-1\leq k\leq t-1$, from (Aletti et al. 2019, Lemma A.5) we get
$[A_{k+1,t}]_{jj}=F_{k+1,t}(\alpha_{j})\,.$
We now prove that $|\boldsymbol{{\rho}}_{t+1}|\stackrel{{\scriptstyle a.s.}}{{\to}}0$. To this end, first notice that $O(|C_{k+1,t}|)=O(|A_{k+1,t}|)$ and, setting $a_{2}^{*}={\mathcal{R}e}(\alpha_{2}^{*})=\phi^{*}-{\mathcal{R}e}(\phi^{*}_{2})$ with $\phi_{2}^{*}$ eigenvalue of $\Phi$, different form $\phi^{*}$, such that ${\mathcal{R}e}(\phi_{2}^{*})=\max_{j}\{{\mathcal{R}e}(\phi_{j})\}$, we have (see (Aletti et al. 2017, Lemma A.4))
@@
and simply $|A_{t+1,t}|=O(1)$ for $k=t$. Moreover, recalling that ${\mathbf{R}}_{A,t+1}=O(t^{-(1+\beta)}){\boldsymbol{1}}$ for some $\beta>0$, we have
@@
Therefore, in all the sequel, without loss of generality, we can assume that ${\mathbf{A}}_{t}$ follows the dynamics (B.7) with ${\mathbf{R}}_{A,t+1}={\boldsymbol{0}}$.
@@
Let $\breve{{\mathbf{A}}}_{t}=V_{\phi}^{\top}{\mathbf{A}}_{t}$ and since ${{\mathbf{A}}}_{\phi,t}=U_{\phi}V_{\phi}^{\top}{\mathbf{A}}_{t}=U_{\phi}\breve{{\mathbf{A}}}_{t}$, it is enough to prove that $\|\breve{{\mathbf{A}}}_{t}\|^{2}$ converges a.s. and in $L^{2}$ to zero. To this end, by multiplying equation (B.7) by $V_{\phi}^{\top}$, we have
@@
(B.14) $E[\Delta{\mathbf{M}}_{t+1}^{\top}Q\Delta{\mathbf{M}}_{t+1}|\mathcal{F}_{t}]$ $=\sum_{j=1}^{N}q_{jj}^{2}E[(\Delta M_{j,t+1})^{2}|\mathcal{F}_{t}]$
@@
which, since $\sum_{t}(\phi^{*}-|\phi|)/(t+1)=+\infty$, by (iv) and (Aletti et al. 2023a, Lemma S1.6), we can conclude that $\|\breve{{\mathbf{A}}}_{t}\|\stackrel{{\scriptstyle a.s./L^{2}}}{{\longrightarrow}}0$, and hence $\breve{{\mathbf{A}}}_{t}\stackrel{{\scriptstyle a.s./L^{2}}}{{\longrightarrow}}{\boldsymbol{0}}$.
@@
From the Frobenius-Perron theory, we know that each eigenvalue with maximum modulus is simple. Let us denote the corresponding right and left eigenvectors by ${\mathbf{v}}_{\phi}$ and ${\mathbf{u}}_{\phi}$ of $\Phi$. Then, set $a_{\phi,t}={\mathbf{v}}_{\phi}^{\top}{\mathbf{A}}_{t}$ so that, since we have ${{\mathbf{A}}}_{\phi,t}={\mathbf{u}}_{\phi}{\mathbf{v}}_{\phi}^{\top}{\mathbf{A}}_{t}={\mathbf{u}}_{\phi}a_{\phi,t}$, it is enough to prove that $|a_{\phi,t}|$ converges to zero almost surely and in $L^{2}$. To this end, by multiplying equation (B.7) by ${\mathbf{v}}_{\phi}^{\top}$, we have
@@
Then, using (B.14), we have that
@@
Since $\sum_{t}(1/(t+1)-\phi^{*}/(t+1)^{2})=+\infty$ and by (iv) and (Aletti et al. 2023a, Lemma S1.6), we can conclude that $|a_{\phi,t}|\stackrel{{\scriptstyle a.s./L^{2}}}{{\longrightarrow}}0$, and hence $a_{\phi,t}\stackrel{{\scriptstyle a.s./L^{2}}}{{\longrightarrow}}0$. ∎∎
B.2. Technical result

### v2

@@ Appendix B Proofs of the main results and technical details for the statistical test
Using (6) and the relation $\zeta_{t+1}(x)/\zeta_{t}(x)=1+O(1/t)$, we obtain
${\mathbf{B}}_{t+1}$ $={\mathbf{B}}_{t}-\frac{1}{t+1}\left(\gamma^{*}{\mathbf{B}}_{t}-{\mathbf{A}}_{t}\right)$ (9)
$\qquad+\frac{1}{t+1}\Delta{\mathbf{M}}_{t+1}+O(\zeta_{t}(1-\gamma^{*})/t^{2}){\mathbf{1}}\,$,
where again $O(\zeta_{t}(1-\gamma^{*})/t^{2})=O(1/t^{1+\gamma^{*}})$. Finally, we observe that, by (5), we have
[…]
and we recall that, as proven before, $\widetilde{A}_{t}={\mathbf{v}}^{\top}{\mathbf{A}}_{t}=\zeta_{t}(1-\gamma^{*}){\mathbf{v}}^{\top}{\mathbf{P}}_{t}\stackrel{{\scriptstyle a.s.}}{{\to}}\widetilde{P}_{\infty}$. Therefore, the pair $({\mathbf{A}}_{t},{\mathbf{B}}_{t})_{t}$ satisfies the dynamics and the conditions required in the general central limit theorem proven in (Aletti, Crimaldi and Ghiglietti 2025, Appendix A), provided we assume ${\mathcal{R}e}(\gamma^{*}_{2})/\gamma^{*}<1/2$, where $\gamma^{*}_{2}$ is an eigenvalue of $\Gamma$ different from $\gamma^{*}$ with highest real part, that is $\gamma^{*}_{2}\in Sp(\Gamma)\setminus\{\gamma^{*}\}$ with ${\mathcal{R}e}(\gamma^{*}_{2})=\max\{{\mathcal{R}e}(\gamma):\gamma\in Sp(\Gamma)\setminus\{\gamma^{*}\}\}$. Hence, we can apply the statistical tools based on that result and described in (Aletti, Crimaldi and Ghiglietti 2025, Appendix B and C). In particular, in the mean-field case we have ${\mathbf{v}}=N^{-1}{\mathbf{1}}$, ${\mathbf{u}}={\mathbf{1}}$ and $\gamma^{*}_{2}=\gamma^{*}(1-\iota)$ so that, when $\iota>1/2$, we have
$(2\iota-1)\frac{\|{\mathbf{S}}_{t}-{\mathbf{v}}^{\top}{\mathbf{S}}_{t}\|^{2}}{{\mathbf{v}}^{\top}{\mathbf{S}}_{t}}=(2\iota-1)\zeta_{t}(\gamma^{*})\frac{\|{\mathbf{B}}_{t}-{\mathbf{v}}^{\top}{\mathbf{B}}_{t}\|^{2}}{{\mathbf{v}}^{\top}{\mathbf{B}}_{t}}\stackrel{{\scriptstyle a.s.}}{{\sim}}(2\iota-1)t^{\gamma^{*}}\frac{\|{\mathbf{B}}_{t}-{\mathbf{v}}^{\top}{\mathbf{B}}_{t}\|^{2}}{{\mathbf{v}}^{\top}{\mathbf{B}}_{t}}\stackrel{{\scriptstyle d}}{{\longrightarrow}}\chi^{2}(N-1).$
B.1 General results
@@
${\mathbf{A}}_{t+1}={\mathbf{A}}_{t}-\frac{1}{t+1}(\phi^{*}I-\Phi^{\top}){\mathbf{A}}_{t}+\frac{1}{t+1}\Phi^{\top}\Delta{{\mathbf{M}}}_{t+1}+{\mathbf{R}}_{A,t+1}\\$ (10)
@@
    ${\mathbf{R}}_{A,t+1}=O(t^{-(1+\beta)}){\mathbf{1}}$ for some $\beta>0$. [^14] The notation $R_{t}=O(s_{t})$ means that $R_{t}$ is a (possibly random) reminder term such that $|R_{t}|\leq Cs_{t}$ for a suitable constant $C$ and for $t$ large enough.
Let $\mathbf{u}$} and $\mathbf{v}$} be the left and the right eigenvectors of $\Phi$ associated to $\phi^{*}$, with strictly positive components and such that ${\mathbf{v}}^{\top}{\mathbf{1}}=1\quad\mbox{and}\quad{\mathbf{v}}^{\top}{\mathbf{u}}=1.$ (Recall that $\phi^{*}$ is real and simple and it is possible to choose the components of these vectors all strictly positive because of the Frobenius-Perron theory). Set $\widetilde{A}_{t}={\mathbf{v}}^{\top}{\mathbf{A}}_{t}$.
@@
By multiplying equation (10) by $\mathbf{v}$$^{\top}$ we obtain
$\widetilde{A}_{t+1}=\widetilde{A}_{t}+\frac{1}{t+1}\phi^{*}\Delta\widetilde{M}_{t+1}+{\mathbf{v}}^{\top}{\mathbf{R}}_{A,t+1}\,$, (11)
@@
$\widetilde{A}_{t+1}-\widetilde{A}_{0}=\sum_{n=0}^{t}(\widetilde{A}_{n+1}-\widetilde{A}_{n})=\phi^{*}\breve{M}_{t+1}+\sum_{n=0}^{t}\widetilde{R}_{A,n+1}\,.$ (12)
Hence, since $\sup_{t}E[\,|\widetilde{A}_{t}|\,]<+\infty$ (by (iii)) and $\sum_{t}1/t^{1+\beta}<+\infty$, we also have $\sup_{t}E[\,|\breve{M}_{t}|\,]<+\infty$. Therefore, $(\breve{M}_{t})_{t}$ is a martingale bounded in $L^{1}$ and so it converges almost surely to an integrable random variable $\breve{M}_{\infty}$. It follows from (12) that $(\widetilde{A}_{t})$ converges almost surely to an integrable random variable $\widetilde{A}_{\infty}$. Moreover, we obtain
$(\widetilde{A}_{t}-\widetilde{A}_{\infty})^{2}$ $=\big(\phi^{*}(\breve{M}_{t}-\breve{M}_{\infty})+\sum_{n\geq t}\widetilde{R}_{A,n+1}\big)^{2}$ (13)
@@
Now, we are going to prove that, under assumption (iv), we have $\sup_{t}(\widetilde{A}_{t})^{2}<+\infty$ so that (by (12)) $(\breve{M}_{t})_{t}$ is a martingale bounded in $L^{2}$ and so $\breve{M}_{\infty}$ is square-integrable and $\breve{M}_{t}$ converges in quadratic mean to it. By (13) this fact obviously implies that $\widetilde{A}_{t}$ converges in quadratic mean to $\widetilde{A}_{\infty}$.
@@
Therefore, from (11), since $\sup_{t}E[\widetilde{A}_{t}]<+\infty$, we get
@@
where we have used (iv) in order to say that the first series is finite. Therefore, assuming ${\mathbf{A}}_{0}$ (and so $\widetilde{A}_{0}$) square-integrable, we have $\sup_{t}E[(\widetilde{A}_{t})^{2}]<+\infty$. ∎
Remark 2 (non-negative case).
Condition (iii) is verified when $\widetilde{A}_{t}$ is non-negative for each $t$. Indeed, for each $t$, we have
@@
and thus, since the last series is finite and $\widetilde{A}_{0}$ is integrable, it follows $\sup_{t}E[\widetilde{A}_{t}]<+\infty$. Hence, if $\widetilde{A}_{t}$ is non-negative, we can conclude that (iii) is verified.
@@
We firstly want to prove that we can neglect the term $R_{A,t+1}$ in the dynamics (10) of ${\mathbf{A}}_{t}$.
@@
$V^{\top}{\mathbf{u}}=U^{\top}{\mathbf{v}}=0,\quad V^{\top}U=U^{\top}V=I,\quad I={\mathbf{u}}{\mathbf{v}}^{\top}+UV^{\top}\,.$ (14)
@@
${\mathbf{A}}_{t+1}$ $=\Big(I-\frac{1}{t+1}U(I\phi^{*}-D)V^{\top}\Big){\mathbf{A}}_{t}$ (15)
@@
${\mathbf{A}}_{t+1}$ $=\ C_{m_{0},t}{\mathbf{A}}_{m_{0}}+\sum_{k=m_{0}}^{t}\frac{1}{k+1}C_{k+1,t}\Phi^{\top}\Delta{\mathbf{M}}_{k+1}$ (16)
@@
Moreover, setting for any $x\in{\mathbb{C}}$ with ${\mathcal{R}e}(x)/(m_{0}+1)<1$, $p_{m_{0}}(x)=1$ and $p_{k}(x)=\prod_{m=m_{0}}^{k}(1-\tfrac{x}{m+1})$ for $k\geq m_{0}$ and $F_{k+1,t}=\frac{p_{t}(x)}{p_{k}(x)}$ for $m_{0}-1\leq k\leq t-1$, from (Aletti, Crimaldi and Ghiglietti 2019, Lemma A.5) we get
$A_{k+1,t}]_{jj}=F_{k+1,t}(\alpha_{j})\,.$
We now prove that $|\boldsymbol{{\rho}}_{t+1}|\stackrel{{\scriptstyle a.s.}}{{\to}}0$. To this end, first notice that $O(|C_{k+1,t}|)=O(|A_{k+1,t}|)$ and, setting $a_{2}^{*}={\mathcal{R}e}(\alpha_{2}^{*})=\phi^{*}-{\mathcal{R}e}(\phi^{*}_{2})$ with $\phi_{2}^{*}$ eigenvalue of $\Phi$, different form $\phi^{*}$, such that ${\mathcal{R}e}(\phi_{2}^{*})=\max_{j}\{{\mathcal{R}e}(\phi_{j})\}$, we have (see (Aletti, Crimaldi and Ghiglietti 2017, Lemma A.4))
@@
and simply $|A_{t+1,t}|=O(1)$ for $k=t$. Moreover, recalling that ${\mathbf{R}}_{A,t+1}=O(t^{-(1+\beta)}){\mathbf{1}}$ for some $\beta>0$, we have
@@
Therefore, in all the sequel, without loss of generality, we can assume that ${\mathbf{A}}_{t}$ follows the dynamics (10) with ${\mathbf{R}}_{A,t+1}={\mathbf{0}}$.
@@
Let $\breve{{\mathbf{A}}}_{t}=V_{\phi}^{\top}{\mathbf{A}}_{t}$ and since ${{\mathbf{A}}}_{\phi,t}=U_{\phi}V_{\phi}^{\top}{\mathbf{A}}_{t}=U_{\phi}\breve{{\mathbf{A}}}_{t}$, it is enough to prove that $\|\breve{{\mathbf{A}}}_{t}\|^{2}$ converges a.s. and in $L^{2}$ to zero. To this end, by multiplying equation (10) by $V_{\phi}^{\top}$, we have
@@
$E[\Delta{\mathbf{M}}_{t+1}^{\top}Q\Delta{\mathbf{M}}_{t+1}|\mathcal{F}_{t}]$ $=\sum_{j=1}^{N}q_{jj}^{2}E[(\Delta M_{j,t+1})^{2}|\mathcal{F}_{t}]$ (17)
@@
which, since $\sum_{t}(\phi^{*}-|\phi|)/(t+1)=+\infty$, by (iv) and (Aletti, Crimaldi and Ghiglietti 2023, Lemma S1.6), we can conclude that $\|\breve{{\mathbf{A}}}_{t}\|\stackrel{{\scriptstyle a.s./L^{2}}}{{\longrightarrow}}0$, and hence $\breve{{\mathbf{A}}}_{t}\stackrel{{\scriptstyle a.s./L^{2}}}{{\longrightarrow}}{\mathbf{0}}$.
@@
From the Frobenius-Perron theory, we know that each eigenvalue with maximum modulus is simple. Let us denote the corresponding right and left eigenvectors by ${\mathbf{v}}_{\phi}$ and ${\mathbf{u}}_{\phi}$ of $\Phi$. Then, set $a_{\phi,t}={\mathbf{v}}_{\phi}^{\top}{\mathbf{A}}_{t}$ so that, since we have ${{\mathbf{A}}}_{\phi,t}={\mathbf{u}}_{\phi}{\mathbf{v}}_{\phi}^{\top}{\mathbf{A}}_{t}={\mathbf{u}}_{\phi}a_{\phi,t}$, it is enough to prove that $|a_{\phi,t}|$ converges to zero almost surely and in $L^{2}$. To this end, by multiplying equation (10) by ${\mathbf{v}}_{\phi}^{\top}$, we have
@@
Then, using (17), we have that
@@
Since $\sum_{t}(1/(t+1)-\phi^{*}/(t+1)^{2})=+\infty$ and by (iv) and (Aletti, Crimaldi and Ghiglietti 2023, Lemma S1.6), we can conclude that $|a_{\phi,t}|\stackrel{{\scriptstyle a.s./L^{2}}}{{\longrightarrow}}0$, and hence $a_{\phi,t}\stackrel{{\scriptstyle a.s./L^{2}}}{{\longrightarrow}}0$. ∎
B.2 Technical result